cometeness and clarity please !
1. Solve the initial value problem using Laplace transforms. ſi ost<5 y - 5y + 4y = 0 t25 y(0) = 0, 7(0) = 1
Answer #1

Consider a differential equation (1 Ost<5 y - 5y +4y=f(t)= y(0)=0, y(0)=1 t25 Apply Laplace transformation on the both sid